OpenStudy (pgxgirl103):

What is the circumference of the circle in terms of pi? The radius is 2.2 and I got 13.82 but the choices are 1.1, 1.21, 2.2 and 4.4 im so confused so could someone please explain what i did wrong?

OpenStudy (bibby):

you're multiplying by pi. it's asking for the answered in terms of pi

OpenStudy (campbell_st):

if the question needs an answer in terms of pi you need to calculate \[C = 2 \times 2.2 \times \pi\]

OpenStudy (pgxgirl103):

what does in terms of pi mean then? I just followed the formula i learned. @bibby

OpenStudy (bibby):

in terms of pi means not multiplying by 3.14 when you see pi. something like 4pi is in terms of pi. 12.45 is not in terms of pi

OpenStudy (pgxgirl103):

So then how do I solve this?

OpenStudy (bibby):

The formula was given to you above \(\large C=2\pi r\) 2*2.2*pi = 4.4 pi and now it's in terms of pi you don't have to simplify further

OpenStudy (pgxgirl103):

So you have to put it in terms of pi? How do you do that?

OpenStudy (bibby):

you just did. 4.4pi is in terms of pi

OpenStudy (pgxgirl103):

So do I not multiply pi?

OpenStudy (bibby):

no, as I stated earlier, if it asks you to keep it in terms of pi, it means it wants pi to be in the answer

OpenStudy (bibby):

so 4.4pi and not 12-13

OpenStudy (pgxgirl103):

I understand now, thanks.